# APRIL: Active Partial Rollouts in Reinforcement Learning to tame long-tail generation

### v1

@@ 3 Experiments
The primary objective of our experiments is to empirically validate the efficacy and robustness of the proposed APRIL. Our investigation is guided by the following key questions:
  • How severe is the long-tail problem in the rollout phase of RL training?
  • Does APRIL significantly improve the efficiency in the rollout (inference) phase of RL?
  • Do the slightly partial off-policy rollouts in APRIL affect convergence and accuracy?
  • Is it compatible with existing RL frameworks and adaptable across hardware platforms?
@@
(a) Rollout and Partial Rollout Configuration
@@
(b) Training Configuration
Figure 3: Distribution of rollout response lengths reveals a pronounced long-tail peak.
3.1 Experimental Setup
Algorithms, Model, and Datasets
@@ 3.1 Experimental Setup
3.2 The Long-Tail Problem in Rollout
To quantify the long-tail problem, we first conducted a pure rollout experiment without model training. In our analysis, we selected representative datasets and generated four rollouts for each input instance, averaging their lengths per instance. A maximum generation length of 32,768 tokens was imposed to fully capture the tail behavior.
As shown in Figure 3, the distribution indicates that more than half of the rollout responses terminate within a few thousand tokens, while a long tail of outliers extends close to the maximum limit. This skewness implies that, in a standard synchronous rollout, the completion time of a batch is dominated by the slowest-generating samples, resulting in substantial GPU idle time, as illustrated in Figure 1.
@@ 3.2 The Long-Tail Problem in Rollout
3.2.1 Performance - Token Throughput
We present a comprehensive evaluation of our active partial rollout (APRIL) represented with the line (—–) against a standard non-partial rollout baseline represented with the line (—–). We measure the rollout throughput, defined as the total number of tokens generated divided by the wall-clock time per rollout iteration. As shown in Figure 4, using the Qwen3-4B model, APRIL consistently improves throughput by 24%, 31%, 35% with GRPO algorithm, and 8%, 11%, 10% with DAPO algorithm, on dapo-math-17k, DeepScaler, and DeepMath-103K datasets. On Qwen3-8B model, the numbers are 26%, 35%, 44% and 8%, 8%, 10% respectively.
@@ 3.2.1 Performance - Token Throughput
3.2.2 Convergence Speed and Accuracy
@@
To investigate this, we tracked the learning curves of models trained with both the baseline and APRIL methods, evaluating their performance on the AIME-2024 benchmark at regular intervals. As illustrated in Figure 5, most of the learning trajectories are nearly identical. Interestingly, APRIL can sometimes achieves a higher final accuracy by 78%). This suggests that incorporating mildly off-policy rollouts enhances rollout diversity, thereby positively influencing both learning dynamics and final model performance. Moreover, training remains stable and can even benefit from the added diversity when updating policies within the RL framework.
We further observed that APRIL provides greater training stability. In several baseline runs, we noted a phenomenon where response lengths abruptly exploded late in training, causing all samples to hit the max_length limit, which in turn led to drops in both rewards and evaluation scores. This issue was not observed in any of our APRIL runs, suggesting that the inclusion of slightly off-policy rollouts may act as a regularizer, preventing the policy from diverging into pathologically long generation modes. Thus, partial rollouts in APRIL may offer the additional benefit of improved training robustness. However, we did not encounter cases where completing a rollout required more than five successive policy versions. If rollouts were to depend on too many earlier policies, such excessive off-policy influence might lead to adverse effects. This remains an open question and a promising direction for future exploration.
3.2.3 Analysis of Percentage of Partial Rollouts and Rollout Length
Figure 6: Proportion of hybrid-policy rollouts in each RL training step.

### v3

@@ 3 Experiments
The objective of our experiments is to validate the effectiveness of APRIL in mitigating long-tail rollout generation by: (1) analyzing the severity of long-tail issue in RL training, especially during rollouts (3.2); (2) evaluating throughput to verify APRIL’s efficiency (3.2.1); (3) assessing convergence and accuracy to verify robustness (3.2.2); and (4) examining potential trade-offs (3.2.3).
3.1 Experimental Setup
Algorithms, Model, and Datasets
Our experiments are conducted on two models Qwen3-4B [^2] https://huggingface.co/Qwen/Qwen3-4B (Yang et al., 2025) and Qwen3-8B [^3] https://huggingface.co/Qwen/Qwen3-8B (Yang et al., 2025) with three math reasoning tasks. We use a diverse set of mathematical reasoning datasets for training, including DAPO-Math-17k  [^4] https://huggingface.co/datasets/open-r1/DAPO-Math-17k-Processed, DeepMath-103K  [^5] https://huggingface.co/datasets/zwhe99/DeepMath-103K, and DeepScaleR  [^6] https://huggingface.co/datasets/agentica-org/DeepScaleR-Preview-Dataset. To evaluate the final performance, we use the AIME-2024 benchmark [^7] http://huggingface.co/datasets/HuggingFaceH4/aime_2024, a collection of recent challenging math reasoning problems. This benchmark serves as a standard for assessing the advanced mathematical reasoning abilities of large language models. To faithfully validate the robustness of the proposed method, we apply APRIL to the two most widely used RL algorithms: GRPO (Mroueh et al., 2025) and DAPO (Yu et al., 2025a).
Hyperparameter Setting and Hardware Platform
[…]
3.2 The Long-Tail Problem in Rollout
Figure 3: Distribution of rollout response lengths reveals a pronounced long-tail peak in rollout of RL.
To quantify the long-tail problem, we first conducted a pure rollout experiment without model training across three datasets. In our analysis, we selected representative datasets and generated four rollouts for each input instance, averaging their lengths per instance. All datasets exhibit long-tail behavior, particularly DeepMath-103K, for which the maximum generation length of 32,768 tokens was imposed to fully capture the long-tail distribution in analysis.
As shown in Figure 3, in all datasets , their rollout length distribution indicates that more than half of the rollout responses terminate within a few thousand tokens, while a long tail of outliers extends close to the maximum limit. This skewness implies that, in a standard synchronous RL framework, the completion time of a batch is dominated by the longest-generating instances. Consequently, substantial GPU idle time occurs during inference, leading to training bubbles, as illustrated in Figure 1 (Above).
3.2.1 Performance - Token Throughput
We present a comprehensive evaluation of our Active Partial Rollout (APRIL) method against a standard non-partial rollout baseline. Specifically, we measure rollout throughput, defined as the total number of tokens generated on 8 GPUs divided by the wall-clock time per rollout iteration. As shown in Table 1, we list the actual throughputs. For the Qwen3-4B model, APRIL consistently improves throughput by 24.4%, 31.8%, 37.7% with GRPO algorithm, and 9.0%, 13.5%, 9.8% with DAPO algorithm across three datasets. As for Qwen3-8B model, the throughputs improve by 26.4%, 34.7%, 49.5% and 8.7%, 8.5%, 10.2%, respectively [^8] Refer to Appendix A.3.1 for a comprehensive throughput comparison across full training trajectories.. Overall average throughput improves by around 22.5%.
Table 1: Comparison of rollout throughput between the baseline (non-partial rollout) and APRIL. APRIL has the higher throughput across datasets: dapo-math-17k , DeepScaler, and DeepMath-103K.
Qwen3-4B Qwen3-8B
Alg. Method
DAPO-
Math-17k
DeepScaleR DeepMath
Avg
Imp.
DAPO-
Math-17k
DeepScaleR DeepMath
Avg
Imp.
GRPO APRIL 12,210 12,183 13,027 31.3% 11,096 10,821 11,594 36.9%
Baseline 9,814 9,252 9,459 8,780 8,031 7,756
DAPO APRIL 13,698 13,916 14,318 10.8% 12,407 12,439 13,075 9.1%
Baseline 12,572 12,257 13,038 11,419 11,470 11,860
3.2.2 Convergence Speed and Accuracy
Table 2: We compare the accuracy of the baseline (non-partial rollout) and APRIL. Overall, APRIL performs comparably to the baseline and, in some cases, achieves slightly better final accuracy (Acc.).
Qwen3-4B Qwen3-8B
Alg. Method
DAPO-
Math-17k
DeepScaleR DeepMath
Avg
Imp.
DAPO-
Math-17k
DeepScaleR DeepMath
Avg
Imp.
GRPO APRIL 65.8% 63.7% 71.2% 3.3% 66.9% 66.0% 68.1% 0.8%
Baseline 65.2% 62.6% 63.0% 66.5% 66.7% 65.5%
DAPO APRIL 68.2% 67.8% 66.3% 4.6% 70.6% 68.9% 71.4% 0.4%
Baseline 68.3% 66.6% 53.5% 69.9% 70.5% 71.8%
A potential concern with APRIL is that introducing off-policy rollouts (i.e., rollouts generated by earlier versions of the policy models , i.e., $\pi(\theta_{k}),\pi(\theta_{k-1}),\ldots,\pi(\theta_{k-m})$, where $m\leq k$, as described in Section 2.2 could destabilize training or degrade convergence and final accuracy in RL training.
To investigate this, we tracked the learning curves of models trained with both the baseline and APRIL methods, evaluating their performance on the AIME-2024 benchmark. As shown in Table 2, for the Qwen3-4B model, APRIL improves accuracy by 0.6%, 1.1%, and 8.2% on GRPO, and by -0.1%, 1.2%, and 12.8% on DAPO. For the Qwen3-8B model, APRIL achieves gains of 0.4%, -0.7%, 2.6% on GRPO, and 0.7%, -1.6%, -0.4% on DAPO. Overall average accuracy improves by around 2.1%.
Overall, APRIL does not degrade performance and, in some cases, achieves notable improvements. This suggests that incorporating mildly off-policy rollouts enhances rollout diversity, thereby positively influencing both learning dynamics and final model performance. Moreover, training remains stable and can even benefit from the added diversity when updating policies within the RL framework. We further observed that APRIL provides greater training stability [^9] Refer to Appendix A.3.2 for detailed accuracy results across training trajectories.. In several baseline runs, we noted a phenomenon where response lengths abruptly exploded late in training, causing all samples to hit the max_length limit, which in turn led to drops in both rewards and evaluation scores. This issue was not observed in any of our APRIL runs, suggesting that the inclusion of slightly off-policy rollouts may act as a regularizer, preventing the policy from diverging into pathologically long generation modes. Thus, partial rollouts in APRIL may offer the additional benefit of improved training robustness. However, we did not encounter any cases where completing a rollout required more than five successive policy versions, i.e., $\pi(\theta_{k}),\pi(\theta_{k-1}),\ldots,\pi(\theta_{k-m})$, where $m\leq 5$. If rollouts were to depend on too many earlier policies, such excessive off-policy influence might lead to adverse effects. This remains an open question and a promising direction for future exploration.
Figure 4: Our analysis indicates that approximately 40% of rollouts incorporate sequences generated by preceding policy models in each RL training step.
3.2.3 Analysis of Percentage of Partial Rollouts and Rollout Length
How much off-policy data does APRIL introduce?
A key feature of our partial rollout method is that tokens from incomplete trajectories are cached and reused in subsequent steps, as illustrated in Figure 2. This process inherently introduces off-policy data. We analyzed the rollout composition of our training batches and found this effect to be substantial. As shown in Figure 4, in a typical step with over-sampling, nearly all unfinished responses among the $256$ additional requests are partially generated and then resumed in the next iteration. Tokens carried over from these truncated instances constitute approximately 40% of the tokens in the subsequent steps. Despite this high proportion, APRIL remains remarkably stable throughout training and even achieves slightly higher final accuracy. This observation suggests that incorporating rollouts from mutiple preceding policy models (within 5 steps as we aforementioned in Section 3.2.2 ) may diversify the rollout data combination and enhance the accuracy in RL training.
In the previous APRIL framework, we focused on partial rollouts at the batch level, where rollout lengths exhibit substantial variability across input instances within a batch, as illustrated in Figure 3. It is important to note that when APRIL determines whether an input instance has completed its rollout, we adopt an instance-level completion criterion. Specifically, in our implementation, each input instance forms a group that encompasses all rollout responses associated with that instance. For example, when n_samples_per_prompt=8=8, each input instance generates eight rollout sequences. The instance, together with its corresponding rollouts, is then collected and passed to the training engine to optimize the policy model. If an instance produces fewer than eight rollouts, the completed rollouts are retained in the data buffer until the remaining rollouts are generated, after which they are collectively used to compute the reward for updating the policy model.
A potential concern is that instance-level granularity might introduce an intra-group long-tail effect, where generated rollouts for a single instance might vary significantly in length. To examine this, we compare rollout-length variability at two levels of granularity: batch level and instance level.
dapo-math-17k DeepScaler DeepMath-103K
GRPO
Refer to caption Refer to caption Refer to caption
DAPO
Refer to caption Refer to caption Refer to caption
(a) $\sigma_{batch-level}$: Batch-level rollout length standard deviation.
dapo-math-17k DeepScaler DeepMath-103K
GRPO
Refer to caption Refer to caption Refer to caption
DAPO
Refer to caption Refer to caption Refer to caption
(b) $\sigma_{instance-level}$: Instance-level rollout length standard deviation.
Figure 5: The x-axis represents the training steps, while the y-axis represents the standard deviation. The baseline (non-partial rollout) is in —–, and APRIL is in —–. We compare the standard deviation of response lengths per iteration at the batch level ($\sigma_{\text{batch-level}}$) and the instance level ($\sigma_{\text{instance-level}}$). The $\sigma_{\text{instance-level}}$ is apparently smaller than $\sigma_{\text{batch-level}}$; thus, there are fewer intra-group long-tail problems.
At the batch level, the standard deviation $\sigma_{batch-level}$ of response lengths across all prompts within an iteration is approximately 4,0004{,}000-4,5004{,}500 tokens across datasets as shown in Figure 5(a). In contrast, at the instance level (within the same input instance), the $\sigma_{instance-level}$ is typically less than 1,5001{,}500 tokens for GRPO algorithm, and 2,4002{,}400 tokens for DAPO algorithm as shown in Figure 5(b). This pronounced gap indicates that responses sampled from the same input instance are considerably more homogeneous in length compared to responses sampled across different instances. Consequently, instance-level completion does not reintroduce a severe intra-group long-tail effect. Due to space limitations, we only present the Qwen-4B results, while more Qwen-8B results are in the appendix [^10] More detailed results about Qwen3-4B and Qwen3-8B are in A.3.3.
3.3 Hardware-Agnostic
To demonstrate both the compatibility and general applicability of APRIL, we integrated it into the widely used open-source RL framework slime and replicated our main experiments across multiple hardware platforms, including NVIDIA H100 and AMD MI300. For deployment on AMD MI300 GPUs, we additionally implemented torch_memory_saver in HIP [^11] AMD HIP is a CUDA-like C++ programming model that enables the same GPU codebase to run efficiently on both AMD (via ROCm) and NVIDIA GPUs. on the AMD side. It is a PyTorch extension designed to alleviate GPU memory pressure during the alternating phases of inference and training under a co-allocated mode, where the training engine and inference engine share the same set of GPUs. This setup requires additional dispatching mechanisms in RL training.
torch_memory_saver introduces a region-based tagging mechanism that allows tensors (e.g., model weights, KV caches, and buffers) to be selectively managed. Specifically, tensors associated with a given tag can be paused, where their physical memory is released while their virtual addresses remain reserved. When needed, these tensors can be resumed by remapping newly allocated physical memory back to the original virtual addresses, thereby preserving CUDA/HIP Graph compatibility. In practice, this mechanism enables selective pausing and multi-stage resumption of memory regions, allowing users to release GPU memory after inference and restore it before subsequent rollouts. The key advantage is that it reduces out-of-memory errors and peak memory consumption without sacrificing performance or requiring modifications to CUDA/HIP Graph execution. With this, it can allow us to conduct RL training in co-allocated mode. The detailed implementation involves extensive discussion of low-level GPU operations and is therefore not emphasized here. The HIP-based implementation [^12] HIP-based implementation torch_memory_saver: https://github.com/fzyzcjy/torch_memory_saver has already been merged into the official upstream GitHub repository and has also been introduced in the online documents131313torch_memory_saver document: https://github.com/sgl-project/sglang/issues/7009, which can be referenced in the footnote below.
4 Related Work
In related work, we provide additional background and analyze the system-level design and algorithmic requirements in existing RL training to further clarify the motivation for creating APRIL.
RL Framework Architectures: Synchronous vs. Asynchronous
A classic mathematical formulation of RL is introduced in Equation 1. In practice, the mainstream implementation follows a synchronous RL paradigm, as exemplified by frameworks such as TRL (von Werra et al., 2020), OpenRLHF (Hu et al., 2025), verl (Sheng et al., 2025), and ROLL (Wang et al., 2025). In this paradigm, the inference and training engines operate as sequential processes. Specifically, when RL training is launched, a batch of instances is dispatched to the inference engine to generate rollouts. All workers simultaneously generate a full batch of experiences using a unified frozen policy model. Only after the entire batch has been collected does the policy update occur in the training engine. This design leverages the inherent on-policy nature of RL, ensuring that the training process consistently utilizes the most recent rollout data. These rollouts, directly sampled from the current policy model, promote training stability and facilitate convergence. However, this simplicity comes at a considerable cost to efficiency. In particular, the system must wait for all instances within a batch to complete rollout generation. Due to the long-tail phenomenon in rollout lengths (Section 3.2), some instances take substantially longer to finish, leading to severe underutilization of computational resources.
In response to the efficiency challenges inherent in synchronous RL, researchers and engineers have increasingly explored more dynamic and decoupled architectures. Fully asynchronous RL frameworks are emerging as a new paradigm in reinforcement learning, distinguished by their ability to overcome the limitations imposed by strictly on-policy rollouts. This architectural evolution builds on earlier advancements in classical methods, notably Asynchronous Advantage Actor-Critic (A3C) (Mnih et al., 2016) and IMPALA (Espeholt et al., 2018), which pioneered similar disaggregated designs. Recent frameworks such as AReaL (Fu et al., 2025), AsyncFlow (Han et al., 2025), StreamRL (Zhong et al., 2025a), and LlamaRL (Wu et al., 2025) have successfully adopted this asynchronous paradigm. More specifically, in an asynchronous RL framework, workers continuously stream rollouts generated by the inference engine into a shared buffer, while the training engine independently retrieves rollouts from this buffer for model updates. The disaggregation of inference and training engines across distinct groups of computational resources enables fine-grained allocation tailored to task-specific demands. This decoupling maximizes hardware utilization and has been shown to deliver substantial throughput improvements. However, it also introduces the challenge of rollout staleness, whereby an increased proportion of off-policy rollouts may be consumed. Such staleness may lead to training instability and, in some cases, to degraded convergence and reduced final accuracy in RL training.
The choice between synchronous and asynchronous paradigms presents a fundamental trade-off between training accuracy and performance. Inspired by this system-level design philosophy, our goal is to preserve the on-policy nature of RL required by many algorithms while simultaneously improving efficiency. To this end, we introduce APRIL, which incorporates an asynchronous mechanism into the rollout generation process to mitigate the long-tail problem and the associated issue of resource underutilization. Crucially, APRIL mitigates long-tail inefficiency without the complexities of full asynchronous decoupling or the burden of managing off-policy staleness. APRIL makes synchronous RL more efficient and robust, narrowing the gap with the performance of asynchronous approaches while preserving the benefits of on-policy rollouts. Our experiments further show that APRIL achieves comparable, and in some cases slightly higher, accuracy than the original synchronous framework.
Rollouts: On-policy vs. Off-policy
The algorithmic foundation of RL is intrinsically linked to the dichotomy of on-policy versus off-policy learning. This fundamental distinction determines how rollout is collected and subsequently used for policy model updates, thereby establishing a critical trade-off between training stability and accuracy versus sample-collection (rollout data) efficiency.
In on-policy learning, PPO (Schulman et al., 2017; Zheng et al., 2023) remains the standard for RL training of LLMs, but it requires fresh rollouts for each update, resulting in sample inefficiency and high computational cost. PPO-based training also involves four distinct LLMs—the policy, reference, reward, and value models—leading to substantial GPU memory overhead. To address these issues, GRPO (Shao et al., 2024; Mroueh et al., 2025) removes the explicit value network by averaging multiple reward-scored responses as a baseline, thereby reducing resource demands. Building on this, DAPO (Yu et al., 2025b) and GSPO (Zheng et al., 2025) extend GRPO to overcome its limitations.
In contrast, off-policy learning updates the policy using rollouts from previous policy versions, alleviating the rollout collection bottleneck but risking instability due to a distribution mismatch between the behavior and target policies. This mismatch can increase gradient variance and harm training stability. Nonetheless, recent work has explored hybrid approaches showing that, with proper management, off-policy data can preserve stability while improving performance. For example, LUFFY (Yan et al., 2025) augments on-policy RL with off-policy reasoning traces, balancing imitation and exploration while applying regularized importance sampling to sustain exploration.  Our method - APRIL’s partial rollout strategy can be viewed as a mild relaxation of strict on-policy training, gaining partial off-policy efficiency benefits while retaining the stability of on-policy RL.
5 Conclusion
In this work, we introduced APRIL, a simple yet effective approach to alleviating long-tail inefficiencies in reinforcement learning for large language models. By reusing partial rollouts and reducing wasted computation, APRIL improves throughput by 22.5% on average across commonly used RL algorithms, accelerates convergence, and, in some cases, achieves an average of 2.1% higher final accuracy across three tasks. In addition, we have made further engineering efforts to support deployment across diverse hardware devices. Beyond these empirical gains, APRIL highlights the importance of addressing both system-level bottlenecks and algorithmic design, offering a practical step toward more efficient and sustainable RL training. We anticipate that its principles will inspire more future research on scalable algorithms and system-level strategies for large-scale RL training.
Ethics Statement
This work focuses on improving the efficiency of reinforcement learning (RL) for large-scale language models by mitigating long-tail inefficiencies in rollout generation. All datasets used in our experiments are publicly available and widely adopted in prior research. Our method, APRIL, is framework- and hardware-agnostic, and is intended to reduce computational overhead, thereby lowering energy consumption and improving sustainability in large-scale RL training. Potential ethical considerations include the broader risks associated with training powerful language models, such as misuse for generating harmful content. While our contribution primarily addresses system-level efficiency, we emphasize that responsible use of LLMs remains critical, and we encourage deployment of our method in alignment with established guidelines for ethical AI research and application.
Reproducibility Statement
We have included our codebase in the supplementary material along with the paper submission. Readers and reviewers can follow the README file to reproduce the results presented in the paper. We also provide the complete hyperparameter settings in the Appendix. Here, we summarize the datasets, algorithms, and hyperparameters for easier review and replication.
Algorithms, Model, and Datasets
(a) Standard (non-partial rollout) and APRIL (partial rollout) configuration
@@ 5 Conclusion
(b) Training Configuration
@@
Appendix A Appendix
A.1 Use of LLMs
We used ChatGPT-5 solely as a general-purpose writing assistant to rephrase sentences and correct grammar and typographical errors in this paper. The model was not involved in research ideation, experimental design, data analysis, or drawing conclusions. The authors take full responsibility for the content of the paper.
A.2 Experimental Setup
Algorithms, Model, and Datasets
(a) Standard (non-partial rollout) and APRIL (partial rollout) configuration
Hyperparameter Value
Rollout Configuration
rollout_batch_size 32
n_samples_per_prompt 8
rollout_max_response_len 16384
rollout_temperature 0.8
n_samples_per_eval_prompt 16
eval_max_response_len 16384
eval_top_p 0.7
Partial Rollout - specific hypermeter
over_sampling_batch_size 64
(b) Training Configuration
Hyperparameter Value
Training Configuration
global_batch_size 256
optimizer Adam
lr 1e-6
weight_decay 0.1
adam_beta1 0.9
adam_beta2 0.98
kl_coef 0.0
eps_clip 0.2
eps_clip_high 0.28
Hardware Platform
Hyperparameter Setting
[…]
A.3 Experiments
A.3.1 Performance - Token Throughput
@@
We present a comprehensive evaluation of our active partial rollout (APRIL) represented with the line (—–) against a standard non-partial rollout baseline represented with the line (—–). We measure the rollout throughput, defined as the total number of tokens generated divided by the wall-clock time per rollout iteration. As shown in Figure 6, we list the actual throughputs. For the Qwen3-4B model, APRIL consistently improves throughput by 24.4%, 31.8%, 37.7% with GRPO algorithm, and 9.0%, 13.5%, 9.8% with DAPO algorithm across three datasets. As for Qwen3-8B model, the throughputs improve by 26.4%, 34.7%, 49.5% and 8.7%, 8.5%, 10.2%, respectively. Overall average throughput improves by around 22.5%.
@@
A.3.2 Convergence Speed and Accuracy
@@
As illustrated in Figure 7, for the Qwen3-4B model, APRIL improves accuracy by 0.6%, 1.1%, and 8.2% on GRPO, and by -0.1%, 1.2%, and 12.8% on DAPO. For the Qwen3-8B model, APRIL achieves gains of 0.4%, -0.7%, 2.6% on GRPO, and 0.7%, -1.6%, -0.4% on DAPO. Overall average accuracy improves by around 2.1%. This suggests that incorporating mildly off-policy rollouts enhances rollout diversity, thereby positively influencing both learning dynamics and final model performance. Moreover, training remains stable and can even benefit from the added diversity when updating policies within the RL framework.
A.3.3 Analysis of Percentage of Partial Rollouts and Rollout Length